OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = −3x + 12 and passes through the point (−1, 6).

OpenStudy (campbell_st):

the equation has the same gradient so its y = -3x + b substitute x = -1 and y = 6 to find the value of b

OpenStudy (anonymous):

6=-3(-1)+b

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

Parallel lines have equal slope.. So, from the equation given find slope that comes out to e -3.. So the other line will have slope = -3 Now use: \[\large y - y_1 = m(x-x_1)\]

OpenStudy (anonymous):

here y1 = 6 and x = -1 and m = -3

OpenStudy (anonymous):

y-6=-3(x+1)

OpenStudy (anonymous):

Absolutely..

OpenStudy (anonymous):

Now distribute -3 inside..

OpenStudy (anonymous):

y-6=-3x-3

OpenStudy (anonymous):

Yes.. add 3x both the sides because you have to bring x and y on left side..

OpenStudy (anonymous):

is the answer y=-3x+3

OpenStudy (anonymous):

Which form in you want to write??

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

Then you are right.... Well Done..

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

can you give the other person a medal to

OpenStudy (anonymous):

Getting how we did it???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I will.. Because he is a person which deserves medal everywhere...

OpenStudy (anonymous):

Welcome dear..

OpenStudy (anonymous):

*who in place of which.. My English is really weak..

OpenStudy (anonymous):

thanx to both of you

OpenStudy (anonymous):

Remember in case when lines are perpendicular then we find the slope of first.. To find the other: remember their slopes are negative reciprocal of each other.. So, if one has a slope of -3, then other will have: \(\large \frac{1}{3}\) Getting?? then use the same concept..

OpenStudy (anonymous):

yes i'd work it the same way right.

OpenStudy (anonymous):

Great.. Take care..